OpenStudy (anonymous):

Solve for the variable: -4h + 2 = 18 A. -4 B. -2 C. 1 D. 6

OpenStudy (anonymous):

a.-4 -4h+2 -2=18-2 -4h=16 h=-4

OpenStudy (anonymous):

thank you so much i was very lost

OpenStudy (anonymous):

h=-4. However try to understand how he cam eo this answer!

OpenStudy (anonymous):

i did he just showed the work

OpenStudy (anonymous):

ok then solve 2x-5=-12

OpenStudy (anonymous):

the more you practice, the better!..